OpenStudy (anonymous):

Which of the following functions is not a linear function? f(x) = -2 f(x) = x - 2 f(x) = 2x f(x) = x 2

OpenStudy (shadowfiend):

Is that last one meant to be \(f(x) = x^2\) ?

OpenStudy (anonymous):

if it is then it's the last one!

OpenStudy (shadowfiend):

Whoa let's not get ahead of ourselves, we'll get to what the answer is ;)

OpenStudy (shadowfiend):

(Hint: it's not the last one.)

OpenStudy (anonymous):

haha

OpenStudy (unklerhaukus):

f(x) = x - 2

OpenStudy (shadowfiend):

^ Also not.

OpenStudy (anonymous):

f(x)=2x

OpenStudy (shadowfiend):

Heh. I actually take my statement back from before, it is \(f(x) = x^2\). Had thought the question was asking something subtly different O:-)

OpenStudy (unklerhaukus):

but how can it be a linear function if it does not cut through the origin?